a boy is 10 years older than his brother. in four years, he will be twice as old as his brother. find the present age of each
Answer by TimothyLamb(4379)

x = boy age
y = brother age
---
x = y + 10
x + 4 = 2(y + 4)
---
put the system of linear equations into standard form
---
x - y = 10
x + 4 = 2y + 8
---
x - y = 10
x - 2y = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = boy age = 16
y = brother age = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php)